Explain generating function.
1
Expert's answer
2021-01-04T14:35:12-0500

The function which provides the method to solve the recurrence relation that is called Generating function.

Let G(t) is a series "G(t)=a_ot^0+a_1t^1+a_2t^2+a_3t^3+......a_nt_n"

Here, "a_o a_1, a_2, a_3...." etc are real number.

So, here G(t) is called generating function of sequence "a_r."

For constant sequence 1,1, 1, 1 ... the generating function is

"G(t)=\frac{1}{1-t}"

So, we can express it as "G(t)=(1-t)^{-1}=1+t+t^2+t^3+......."